# Unraveling and Mitigating Retriever Inconsistencies in Retrieval-Augmented Large Language Models

### v1

@@ Abstract
Although Retrieval-Augmented Large Language Models (RALMs) demonstrate their superiority in terms of factuality, they do not consistently outperform the original retrieval-free Language Models (LMs). Our experiments reveal that this example-level performance inconsistency exists not only between retrieval-augmented and retrieval-free LM but also among different retrievers. To understand this phenomenon, we investigate the degeneration behavior of RALMs and theoretically decompose it into four categories. Further analysis based on our decomposition reveals that the innate difference in knowledge sources and the unpredictable degeneration of the reader model contribute most to the inconsistency. Drawing from our analysis, we introduce Ensemble of Retrievers (EoR), a trainable framework that can adaptively retrieve from different knowledge sources and effectively decrease unpredictable reader errors. Our experiments on Open Domain Question Answering show that EoR substantially improves performance over the RALM with a single retriever by considerably reducing inconsistent behaviors.
@@
Mingda Li1 Xinyu Li1 Yifan Chen1 Wenfeng Xuan2 Weinan Zhang1 1Research Center for Social Computing and Information Retrieval Harbin Institute of Technology, China 2XVERSE Technology Inc., China {mdli, xyli, yfchen, wnzhang}@ir.hit.edu.cn {johnxuan}@xverse.cn
@@ 2.1 Experimental Setup
We adopt the zero-shot in-context RALM (Ram et al., 2023) which directly prepends the retrieved documents to the input query based on the prompt template (see Appendix C.2). This naive yet efficient framework have been widely used in recent works (Gao et al., 2023b). We employ Llama2-chat7B, 13B (Touvron et al., 2023) and ChatGPT222gpt-3.5-turbo-instruct as our base LM and perform greedy-search on all response generations to reduce the hallucination brought by sampling and guarantee reproducibility.
@@
Evaluation Metrics: We need two kinds of metrics, one to evaluate the correctness of answers and one to evaluate the example-level inconsistency. Following Kamalloo et al. (2023), we adopt BEM score Bulian et al. (2022), a semantic similarity metric specifically developed for QA tasks, to evaluate QA accuracy with threshold 0.8.777We also tried Exact Match, the results are similar. It is reported to have good correlation with humans and cope with syntactical variation of answers.
@@ 4.3 Experimental Setting
Same as section 2.1, we use NQ, WebQ and TriviaQA as our experiment datasets and Llama2-chat7B, 13B and ChatGPT as our base LM. We search parameters on the validation split (train split for WebQ) and report performance on the test split. For ChatGPT, we randomly sample 500 questions from each split same as Section 2.1. We use BEM accuracy, MRWR and MRLR as evaluation metrics.
@@ 7 Acknowledgements
We thank the HIT SCIR-DT group members for their valuable discussions and insightful feed-back. This research was supported by the National Key Research and Development Program (No. 2022YFF0902100) and Du Xiaoman (Beijing) Science Technology Co., Ltd.
@@ Appendix D Complete Figures
Appendix E Additional Experiments
Base $\mathcal{R}$ NQ WebQ TriviaQA
Models BEM EM F1 BEM EM F1 BEM EM F1
Llama2-chat7B ReFree 34.35 26.70 23.96 51.82 38.34 35.10 55.57 51.29 52.48
Top $\mathcal{R}$ 55.57 46.32 42.24 56.25 43.16 37.49 77.34 72.59 73.18
EoR 58.92 50.22 45.68 59.45 46.16 40.95 80.62 75.85 77.24
Llama2-chat13B ReFree 46.43 35.84 20.23 58.76 44.88 27.94 66.36 60.54 48.42
Top $\mathcal{R}$ 62.30 50.94 46.18 62.20 49.11 37.64 82.37 75.99 72.29
EoR 64.24 53.07 38.69 60.63 47.19 28.59 83.80 77.77 63.27
ChatGPT ReFree 52.40 44.60 41.82 59.20 46.00 39.91 81.60 76.8 77.76
Top $\mathcal{R}$ 60.20 49.6 46.20 60.60 49.2 38.55 84.40 80.6 79.09
EoR 63.00 52.80 50.80 61.60 50.60 39.86 87.40 83.00 82.75

### v5

@@ Abstract
Although Retrieval-Augmented Large Language Models (RALMs) demonstrate their superiority in terms of factuality, they do not consistently outperform the original retrieval-free Language Models (LMs). Our experiments reveal that this example-level performance inconsistency exists not only between retrieval-augmented and retrieval-free LM but also among different retrievers. To understand this phenomenon, we investigate the degeneration behavior of RALMs and theoretically decompose it into four categories. Further analysis based on our decomposition reveals that the innate difference in knowledge sources and the unpredictable degeneration of the reader model contribute most to the inconsistency. Drawing from our analysis, we introduce Ensemble of Retrievers (EoR), a trainable framework that can adaptively retrieve from different knowledge sources and effectively decrease unpredictable reader errors. Our experiments on Open Domain Question Answering show that EoR substantially improves performance over the RALM with a single retriever by considerably reducing inconsistent behaviors. [^1] Our code and data are available at https://github.com/mingdali6717/Ensemble-of-Retrievers
@@
Mingda Li1 Xinyu Li1 Yifan Chen1 Wenfeng Xuan2 Weinan Zhang1thanks: Corresponding author 1Research Center for Social Computing and Information Retrieval Harbin Institute of Technology, China 2XVERSE Technology Inc., China {mdli, xyli, yfchen, wnzhang}@ir.hit.edu.cn {johnxuan}@xverse.cn
@@ 2.1 Experimental Setup
We adopt the zero-shot in-context RALM (Ram et al., 2023) which directly prepends the retrieved documents to the input query based on the prompt template (see Appendix C.2). This naive yet efficient framework have been widely used in recent works (Gao et al., 2023b). We employ Llama2-chat7B, 13B (Touvron et al., 2023) and ChatGPT333gpt-3.5-turbo-instruct as our base LM and perform greedy-search on all response generations to reduce the hallucination brought by sampling and guarantee reproducibility.
@@
Evaluation Metrics: We need two kinds of metrics, one to evaluate the correctness of answers and one to evaluate the example-level inconsistency. Following Kamalloo et al. (2023), we adopt BEM score Bulian et al. (2022), a semantic similarity metric specifically developed for QA tasks, to evaluate QA accuracy with threshold 0.8.888We also tried Exact Match, the results are similar. It is reported to have good correlation with humans and cope with syntactical variation of answers.
@@ 4.3 Experimental Setting
Same as section 2.1, we use NQ, WebQ and TriviaQA as our experiment datasets and Llama2-chat7B, 13B and ChatGPT as our base LM. We search parameters on the validation split (train split for WebQ) and report performance on the test split. For ChatGPT, we randomly sample 500 questions from each split same as Section 2.1. We use BEM accuracy, EM and MRLR as evaluation metrics.
@@ 7 Acknowledgements
We thank the HIT SCIR-DT group members for their valuable discussions and insightful feedback. This research was supported by the National Key Research and Development Program (No. 2022YFF0902100), Du Xiaoman (Beijing) Science Technology Co., Ltd and Nature Scientific Foundation of Heilongjiang Province (YQ2021F006).